OpenStudy (anonymous):

A jar contains 5 peanut butter cookies, 3 carmel delight cookies, and 2 lemon cookies. If two cookies are selected in succession, find the probability of of selecting a peanut butter followed by carmel, if you replace the first; without replacing the first

OpenStudy (anonymous):

If you replace the first , probability = 5/10 * 3/10 = 15/100 = 3/20 If you don't replace it, probability= 5/10 * 3/9 = 15/90 = 1/6

OpenStudy (anonymous):

Thank you :)